OpenStudy (anonymous):

Write the point-slope form of the equation passing through (5, -1) with a slope of 6.

OpenStudy (anonymous):

Point slope form is \[y-y_1=m(x-x_1)\]so just plug in your values to that equation

OpenStudy (anonymous):

the answer is y-(-1)=6(X-5)